# Self-organizing railway traffic management

### v1

@@ VIII Conclusions and future work
In a thorough experimental analysis covering a portion of the Italian network shared by different railway undertakings operating mixed traffic, we showed that the traffic self-organization was able to outperform the state-of-the-art centralized decision-making algorithm used as a benchmark. This result was achieved thanks to the careful design of the various modules of the self-organization process. These modules allow the suitable instance decomposition, to profit from the increased problem tractability without paying the price of making myopic decisions.
In future research, we will deeply investigate the impact of each module configuration on the self-organization performance. Moreover, we will propose alternative designs for some of them, relaying for example on machine learning techniques to further improve the instance decomposition.

### v2

@@ VIII Conclusions and future work
In a thorough experimental analysis covering a portion of the Italian network shared by different railway undertakings operating mixed traffic, we showed that the traffic self-organization often found solutions that are very close to the optimal ones, and was able to outperform the state-of-the-art centralized decision-making algorithm used as a benchmark. This result was achieved thanks to the careful design of the various modules of the self-organization process. These modules allow the suitable instance decomposition, to profit from the increased problem tractability without paying the price of making myopic decisions.
In future research, we will deeply investigate the impact of each module configuration on the self-organization performance, also linked to different types and sources of traffic perturbations. Moreover, we will propose alternative designs for some of them, relying for example on machine learning techniques to further improve the instance decomposition,
and we will exploit graph theoretical analysis to try to explain the relations between modules’ configuration and design, and system performance. The impact of different traffic characteristics also deserves future studies, in particular to understand the effect of having a few competing companies, each operating many trains.